# Comments, corrections, and related references welcomed, as always! ed August 22, 2025

### v2

@@ Abstract.
The latter version may be revised more frequently than the former. After publication, any updates, errata, related references, etc., found will be noted at http://math.berkeley.edu/~gbergman/papers/abstracts .
1. Note on terminology.
@@
2. The category of nonempty semigroups has no prime objects.
@@
(iv) Given a semigroup $S,$ a skeleton $S^{\prime}$ of $S$ will mean any subsemigroup of $S$ which consists of all the product elements, and one and only one representative of each action equivalence class that contains no product elements.
@@
If (i) holds, let us assume without loss of generality that $S=S_{0}\times\textup{Null}(\kappa).$ Note that for each $s\in S_{0},$ the $\kappa\!+\!1$ elements $(s,x)$ $(x\in\textup{Null}(\kappa))$ are action equivalent, and at most one (the one with $x$ the product element of $\textup{Null}(\kappa))$ can be a product element, so taking the union of these families over the elements $x$ of an action equivalence class of $S_{0},$ and writing $A_{0}$ for the set of $(s,x)$ in this union with $x$ the product element of $\textup{Null}(\kappa),$ and $A_{1}$ for the set of $(s,x)$ with non-product $x,$ we get the desired decomposition of the induced action equivalence class of $S.$
@@
Since all 3\!3\!-fold products equal $z,$ $S$ is associative. Note that the action equivalence classes of $S$ are, on the one hand, the sets
@@
and one more equivalence class, comprising all the remaining elements:
@@
The above result does not, however, apply with $C$ taken to be the category of all finite semigroups, since the condition that $C$ contain all semigroups $S$ with zero element is not limited to finite $S.$ Indeed, the proof of Lemma 2.4 used semigroups $\textup{Null}(\kappa)$ in a way that required $\kappa$ to be infinite even when the given semigroup $S$ was finite. Perhaps there is some variant argument that would use only finite semigroups in such cases. Not knowing whether there is, I ask
@@
3. Preparation for some positive results.
@@
Some of the preliminary results we will develop to help us prove these conclusions are true in more general contexts than those in which we will use them. In particular, though we will eventually use the next lemma with $G$ the additive group of integers, the proof does not require commutativity, so I will use multiplicative notation till we specialize further.
The referee indicated that that lemma was probably known. The one published result I am now aware of that implies it is [6, Lemma 3.1], where the authors simply say it holds by “easy computation”. The proof below, suggested by Benjamin Steinberg, is neater and shorter than my original version.
@@
(The analog of the above lemma with the group $G$ replaced by a semigroup, or even a monoid, is not true. For instance, for any $n>1,$ let $S$ be the additive semigroup of all positive integers, $A$ the subsemigroup of integers $\geq n,$ $G$ the additive monoid $A\cup\{0\},$ and $\pi:A\to G$ the inclusion map. Likewise, the analog of the lemma with the assumption that $A$ be an ideal of $S$ weakened to say that it is a left or right ideal fails. For instance, if $S$ is the free semigroup on two elements $s$ and $t,$ and we take for $A$ the left ideal generated by $t,$ then we find that $A$ is free as a semigroup on the set $\{s^{i}\,t\,|\,i\geq 0\},$ so for any nontrivial group $G$ there exist homomorphisms $\pi:A\to G$ acting arbitrarily on these generators. But such a homomorphism not satisfying the relations $\pi(s^{i}\,t)\,\pi(t)^{-1}\pi(s^{j}t)=\pi(s^{i+j}t)$ for all natural numbers $i$ and $j$ cannot be extended to $S.)$
@@
Apply Corollary 3.2 with $I=\{0,1\},$ $S_{0}=S,$ $S_{1}=T,$ and for $s\in S,$ $t\in T,$ define $\pi_{S}(s)=\hat{\pi}(s,e)$ and $\pi_{T}(t)=\hat{\pi}(e,t).$ Then (3.5) and the mutual centralization assertion both follow from the fact that $\hat{\pi}$ respects multiplication. Uniqueness can be seen by noting that any $\pi_{S}$ and $\pi_{T}$ with the indicated properties extend to homomorphisms $S\cup\{e\}\to G$ and $T\cup\{e\}\to G,$ and calling on the uniqueness condition of Corollary 3.2. ∎
@@
If, in the context of Lemma 3.1, $G$ is an additive subgroup of the real numbers, and $\pi:A\to G$ assumes only nonnegative values, then $\hat{\pi}:S\to G$ also assumes only nonnegative values.
@@
Given $S,~\pi,$ and $\nu$ as above, note that (3.7) and (3.8) together imply that for every $i\geq 0,$ we have the same statements with $\nu$ replaced by $\nu^{i}$ and “+1+1” replaced by “+i+i”; and likewise for (3.9).
@@
Since $\pi$ is a semigroup homomorphism, its value on a product of two elements of $U$ will be $2k,$ hence by the preceding observation with $i=k,$ we can use the $\nu^{k}\!-$version” of (3.8) to define a binary operation on $U$:
@@
This proves the main assertion of the lemma, which is the “if” direction of the final sentence. The forward implication of that final sentence is straightforward. ∎
@@
Note that the right-hand side does not depend on $t^{\prime},$ hence neither does the left-hand side; and by the right-left dual argument, it likewise does not depend on $t,$ so we get the same value on replacing the subscript $t\,t^{\prime}$ by any element of $T^{\prime}.$ Here the argument of $\nu,$ a product of two arbitrary element of $S,$ can be any element of $S^{\prime}.$ This proves the first line of (3.16), allowing us to adopt the notation in the first line of (3.17). The first equality of the first line of (3.19) follows; and by interchanging right and left multiplication, and the roles of $S$ and $T,$ as appropriate, we get (3.18) and the remaining equalities of (3.16) and (3.19). ∎
@@
4. Primeness results, at last.
We are at last ready to note some subcategories of $\mathbf{Semigp}$ in which we can use the above results to prove some or all of the semigroups $\mathbb{N}\!+\!k$ prime. These categories cannot, of course, contain the semigroups $\textup{Null}(\kappa)$ that were used to prove non -primeness in §2. Here are some restrictions that suffice.
@@
So suppose, rather, that for some $s\in S,$ $t,\,t^{\prime}\in T$ we had $\nu_{t}(s)\neq\nu_{t^{\prime}}(s).$ We now invoke weak cancellativity. By left-right symmetry, we may assume without loss of generality that there is some $s^{\prime}\in S$ such that $\nu_{t}(s)\,s^{\prime}\neq\nu_{t^{\prime}}(s)\,s^{\prime}.$ But by (3.19), both sides equal $\nu(s\,s^{\prime}),$ a contradiction.
@@
The context of that result concerns categories of algebras whose operations include a binary operation denoted ++ (but not required to be commutative or even associative) and a zeroary operation $0,$ such that $\{0\}$ is a subalgebra, and the identity $0+x=x=x+0$ holds [9, start of §5.4]. The result also involves the “center” of such an algebra, a complicated concept developed in [9, §5.5]; but all we will need to know about it below that for all elements $b$ in the center, there exists an element $c$ such that $b+c=0$ [9, Definitions 5.11 and 5.12]. (Warning: the word “center” is used in different, though overlapping ways in other chapters of [9].) The result is
@@
Note that primeness in such a category implies primeness in any subvariety, and that a monoid with no nontrivial invertible elements (or more generally, with only finitely many) necessarily satisfies the “finite center” condition of (4.5). Some examples in the variety of monoids that (4.5) shows to be prime include:  (i) all nontrivial submonoids of $\mathbb{N},$ (ii) all submonoids $S$ of $\mathbb{N}\times\mathbb{N}$ such that the submonoids $S\cap(\mathbb{N}\times\{0\})$ and $S\cap(\{0\}\times\mathbb{N})$ are both nontrivial, and $S$ is strictly larger than their sum (e.g., $\{(a,\,b)\in\mathbb{N}\times\mathbb{N}\ |\ a\equiv b\ (\mathrm{mod}~n)\}$ for any $n>1,$ or $\mathbb{N}\times\mathbb{N}\setminus\{(1,0)\}),$ (iii) all submonoids $S$ of $\mathbb{N}\times\mathbb{N}$ such that $S\cap(\mathbb{N}\times\{0\})=\{(0,0)\},$ while $S\cap(\mathbb{N}\times\{1\})$ and $S\cap(\{0\}\times\mathbb{N})$ each have more than one element, and (iv) all free monoids on more than one (possibly infinitely many) generators. (That none of these examples is a nontrivial directly product takes a bit of thought in case (ii) and more in case (iii). I leave these as exercises for the reader.)
@@
On the other hand, we shall see in Example 6.2 below that the additive group of integers is not prime (in the language of that section, not “Tarski-prime”) in the category of groups, from which it easily follows that it is also non-prime in the categories of semigroups and monoids considered above.
5. Two other questions from [9].
@@
An algebra of type <1,1><\!1,1\!> can be thought of as a set given with an action on it of the free semigroup on two generators; so in that way the question has a somewhat similar flavor to those studied in the preceding sections.
@@ 6. Some results and examples for groups, comparing
Tarski-primeness and the two sorts of Rhodes-primeness.
@@
(iv)  we shall call $G$ Rhodes-prime with respect to direct products if, whenever $G$ is a subquotient of a group in the category in question, and the latter group can be written as a direct product $G_{0}\times G_{1},$ then $G$ is a subquotient of one of $G_{0}$ or $G_{1}$ (condition introduced at [11, p. 482, line 6]), and similarly,
@@
(What we name a subquotient of $G$ in (ii) above was earlier called a factor of $G,$ e.g., [10, p. 1 et seq.], is called a divisor of $G$ in [11, Def. 1.2.25, p. 32], and is now often called a section of $G$ by group theorists.)
We will also follow the common conventions of generally writing $e$ for the identity elements of groups, though for particular groups we may use notation appropriate for them, e.g., 0 for the identity element of the additive group $\mathbb{Z},$ and of calling a group $G$ nontrivial if it has more than one element. (We use “nontrivial” in that sense in the line introducing (iii)-(v) above. In [11] the trivial semigroup is counted as prime with respect to direct and semidirect products, but here we have excluded that group to make the comparison between the Rhodes-primeness conditions and that of Tarski-primeness more straightforward.)
@@
Note that neither of the groups $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ and $\mathbb{Z}_{q}\rtimes\mathbb{Z}$ has a direct product decomposition with a factor isomorphic to $\mathbb{Z}.$ E.g., if $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ did, then a generator of that factor would be a central element of infinite order, hence would have, under the given description of the group, the form (0,ma)(0,\,ma) $(m\neq 0).$ But that would make it an a\!a\!-th power in the whole group, while looking at the $\mathbb{Z}\!$-component in such a direct product decomposition, we see that it can’t be an a\!a\!-th power for any $a>1.$ (The groups $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ and $\mathbb{Z}_{q}\rtimes\mathbb{Z}$ are in fact directly indecomposable, but the above argument gives as much as we need.)
On the other hand, suppose we rewrite $G$ as $(\mathbb{Z}_{p}\times\mathbb{Z}_{q})\rtimes(\mathbb{Z}\times\mathbb{Z}),$ using the action of $\mathbb{Z}\times\mathbb{Z}$ under which the first factor $\mathbb{Z}$ affects only the first factor of $\mathbb{Z}_{p}\times\mathbb{Z}_{q},$ and the second only affects the second.
Since $a$ and $b$ are relatively prime, the subgroup of $\mathbb{Z}\times\mathbb{Z}$ generated by the element $(a,b)$ is a direct factor thereof. Explicitly, we can find integers $c$ and $d$ such that bcad=1,bc-ad=1, and use the invertible matrix $(\begin{smallmatrix}c&d\\ a&b\end{smallmatrix})$ to convert the given direct product decomposition of $\mathbb{Z}\times\mathbb{Z}$ into one in which $(a,b)$ generates the second factor. (That matrix or its inverse necessarily has some negative entries, which can be thought of as the way the fact that $\mathbb{Z}$ includes negative integers comes in.) By (6.2), that second factor acts trivially on $\mathbb{Z}_{p}\times\mathbb{Z}_{q}.$
@@
For finite groups, the relationship between these primeness conditions is quite different. A key result is the Krull-Schmidt Theorem [5, Theorem 3.8, p. 86], which says that every group with ACC and DCC on normal subgroups has a factorization as a direct product of finitely many nontrivial groups which are not themselves nontrivial direct products, and that the factors in such a decomposition are unique up to rearrangement and isomorphism. Thus, every group with those chain conditions which is not a nontrivial direct product is Tarski-prime in the category of such groups. In particular, every nontrivial finite group which is not a nontrivial direct product is Tarski-prime in the category of finite groups. (For the corresponding result about finite algebras much more general than groups, of which this is a special case, see (4.5) above, together with the paragraph that precedes that display and the second paragraph following it.)
Recall that if a group $G$ has a unique minimal nontrivial normal subgroup $M,$ then $G$ is called monolithic, and $M$ is called the monolith of $G$ [10, sentence beginning at bottom of p. 146].
We can now prove
@@
Though $N$ is a minimal nontrivial normal subgroup of $G,$ it need not be simple, but it will be characteristically simple, i.e., it will have no proper nontrivial subgroups invariant under all its automorphisms (since these include conjugation by all members of $G).$ Hence by [12, 3.3.15, p. 87, and sentence at end of proof, p. 88], $N$ is a direct product of (mutually isomorphic) simple groups. Hence $N^{|G/N|}$ is a direct product of those same simple groups, with possibly more repetitions.
@@
Given a finite group $G$ that is not Rhodes-prime with respect to direct products, we shall show that neither (i) nor (ii) can hold. The proof that (iii) implies Rhodes-primeness will be more straightforward.
Assuming non-Rhodes-primeness with respect to direct products, let us write $G$ as a homomorphic image $f(H)$ of a subgroup $H$ of a direct product $G_{0}\times G_{1}$ of finite groups, such that $G$ is not a subquotient of $G_{0}$ or of $G_{1}.$ We can clearly replace $G_{0}$ and $G_{1}$ by the subgroups given by the projections of $H$ onto those two factors, so assume that those projections are surjective; i.e., that $H$ is a subdirect product of $G_{0}$ and $G_{1}.$ Let us further assume that, for the given group $G,$ the groups $G_{0},$ $G_{1},$ and $H$ are chosen so as to minimize the order of $H.$
@@
Now let $M_{0}\times\{e\}$ be a nontrivial subgroup of $H_{0}\times\{e\}$ minimal for being normal in $H.$ Since $M_{0}\subseteq H_{0},$ (6.5) tells us that $M_{0}\times\{e\}$ maps one-to-one into $G,$ hence its image is a minimal normal subgroup of $G;$ hence that image is $M.$ Combining with the corresponding observation about $\{e\}\times H_{1},$ we get
@@
Next, suppose as in (ii) that $G$ is a semidirect product $M\rtimes K$ for some $K<G.$ Since $K$ has trivial intersection with the monolith $M$ of $G,$
@@
Returning to what we proved earlier about $H,$ note that if $f^{-1}(K)<H$ had nontrivial intersection with $H_{0},$ then this would centralize $H_{1},$ hence its image in $G$ would be a subgroup of $K,$ nontrivial by (6.5), that centralized the image of $H_{1},$ which we saw contains $M,$ contradicting (6.8). So $f^{-1}(K)$ has trivial intersection with $H_{0};$ and similarly with $H_{1}.$ In view of (6.4), this forces $f^{-1}(K)$ to be the graph of an isomorphism between subgroups $K_{0}<G_{0}$ and $K_{1}<G_{1},$ each isomorphic to $K,$ and we see that these act on $M_{0}$ and $M_{1}$ as $K$ acts on $M.$ Hence the subgroup $M_{0}\,K_{0}<G_{0}$ (and likewise $M_{1}\,K_{1}<G_{1})$ is isomorphic to $MK=G,$ again contradicting our assumption that $G$ is not a subquotient of $G_{0}$ or $G_{1}.$ This gives the desired contradiction to (ii) for $G$ not Rhodes-prime with respect to direct products.
To prove Rhodes-primeness with respect to direct products in case (iii), suppose a subgroup $H$ of a finite group $G_{0}\times G_{1}$ maps surjectively to $G=\mathbb{Z}_{p^{n}}.$ Then an element mapping to a generator of $G$ must have order divisible by $p^{n}.$ Since the order of an element of $G_{0}\times G_{1}$ is the least common multiple of the orders of its components, one of those components must have order divisible by $p^{n},$ hence some power of that element will have order exactly $p^{n},$ hence the subgroup of $G_{0}$ or $G_{1}$ that it generates will have that order, making $G=\mathbb{Z}_{p^{n}}$ a subquotient of $G_{0}$ or $G_{1}.$

### v3

# Comments, corrections, and related references welcomed, as always! ed October 26, 2025

@@ Abstract.
After publication, any updates, errata, related references, etc., found will be noted at http://math.berkeley.edu/~gbergman/papers/abstracts .
1. Note on terminology
@@
2. The category of nonempty semigroups has no prime objects
@@
(iv) Given a semigroup $S,$ a skeleton $S^{\prime}$ of $S$ will mean any subsemigroup of $S$ which consists of all the product elements, and exactly one representative of each action equivalence class that contains no product elements.
@@
If (i) holds, let us assume without loss of generality that $S=S_{0}\times\textup{Null}(\kappa).$ Note that for each $s\in S_{0},$ the $\kappa\!+\!1$ elements $(s,x)$ $(x\in\textup{Null}(\kappa))$ are action equivalent, and at most one (the one with $x$ the product element of $\textup{Null}(\kappa))$ can be a product element, so taking the union of these families over the elements $x$ of an action equivalence class of $S_{0},$ and writing $A_{0}$ for the set of $(s,x)$ in this union with $x$ the product element of $\textup{Null}(\kappa),$ and $A_{1}$ for the set of $(s,x)$ with non-product $x,$ we get a decomposition of the induced action equivalence class of $S$ as in (ii).
@@
Since all 3\!3\!-fold products equal $z,$ $S$ is associative. Note that the action equivalence classes of $S$ are of two sorts: on the one hand, the sets
@@
and one more equivalence class, comprising the remaining elements:
@@
The above result does not, however, apply with $C$ taken to be the category of all finite semigroups, since the condition that $C$ contain all semigroups $S$ with zero element equal to all three-fold products is not limited to finite $S.$ Indeed, the proof of Lemma 2.4 used semigroups $\textup{Null}(\kappa)$ in a way that required $\kappa$ to be infinite even when the given semigroup $S$ was finite. Perhaps there is some variant argument that would use only finite semigroups in such cases. Not knowing whether there is, I ask
@@
3. Preparation for some positive results
@@
Some of the preliminary results we will develop to help us show this are true in more general contexts than those in which we will use them. In particular, though we will eventually use the next lemma with $G$ the additive group of integers, the proof of the lemma does not require $G$ to be commutative, so I will use multiplicative notation till we specialize further.
The referee indicated that that lemma was probably known. The one published result I am now aware of that implies it is [6, Lemma 3.1], where the authors say it holds by “easy computation” but give no further details.. The proof below, suggested by Benjamin Steinberg, is neater and shorter than my original version.
@@
(The analog of the above lemma with the group $G$ replaced by a semigroup, or even a monoid, is not true. For instance, for any $n\geq 1,$ let $S$ be the additive semigroup of all positive integers, $A$ the subsemigroup of integers >n,>n, $G$ the additive monoid $A\cup\{0\},$ and $\pi:A\to G$ the inclusion map. Likewise, the analog of the lemma with the assumption that $A$ be an ideal of $S$ weakened to say that it is a left or right ideal fails. For instance, if $S$ is the free semigroup on two elements $s$ and $t,$ and we take for $A$ the left ideal generated by $t,$ then we find that $A$ is free as a semigroup on the set $\{s^{i}\,t\,|\,i\geq 0\},$ so for any nontrivial group $G$ there exist homomorphisms $\pi:A\to G$ acting arbitrarily on these generators. But such a homomorphism not satisfying the relations $\pi(s^{i}\,t)\,\pi(t)^{-1}\pi(s^{j}t)=\pi(s^{i+j}t)$ for all natural numbers $i$ and $j$ cannot be extended to $S.)$
@@
Apply Corollary 3.2 with $I=\{0,1\},$ $S_{0}=S,$ $S_{1}=T,$ and for $s\in S,$ $t\in T,$ define $\pi_{S}(s)=\hat{\pi}(s,e)$ and $\pi_{T}(t)=\hat{\pi}(e,t).$ From the fact that $\hat{\pi}$ respects multiplication, (3.5) and the mutual centralization assertion both follow. Uniqueness can be seen by noting that any $\pi_{S}$ and $\pi_{T}$ with the indicated properties extend to homomorphisms $S\cup\{e\}\to G$ and $T\cup\{e\}\to G,$ and calling on the uniqueness condition of Corollary 3.2. ∎
@@
In the context of Lemma 3.1, if $G$ is an additive subgroup of the real numbers, and $\pi:A\to G$ assumes only nonnegative values, then $\hat{\pi}:S\to G$ also assumes only nonnegative values.
@@
Given $S,~\pi,$ and $\nu$ as above, note that (3.7) and (3.8) together imply that for every $i\geq 0,$ we have the same statements with $\nu$ replaced by $\nu^{i}$ and “+1+1” replaced by “+i+i”; and that (3.9) likewise holds with $\nu$ replaced by $\nu^{i}.$
@@
Since $\pi$ is a semigroup homomorphism, its value on a product of two elements of $U$ will be $2k,$ hence by the $\nu^{k}\!-$version” of (3.8) we can define a binary operation on $U$:
@@
This proves the main assertion of the lemma, namely the “if” direction of the final sentence. The forward implication of that final sentence is straightforward. ∎
@@
Note that the right-hand side does not depend on $t^{\prime},$ hence neither does the left-hand side; and by the right-left dual argument, it likewise does not depend on $t,$ so we get the same value on replacing the subscript $t\,t^{\prime}$ by any element of $T^{\prime}.$ Here the argument of $\nu,$ a product of two arbitrary elements of $S,$ can be any element of $S^{\prime}.$ This proves the first line of (3.16), allowing us to adopt the notation in the first line of (3.17). The first equality of the first line of (3.19) follows; and by interchanging right and left multiplication, and the roles of $S$ and $T,$ as appropriate, we get the remaining equalities of (3.16) and (3.19). ∎
@@
4. Primeness results, at last
We are now ready to note some subcategories of $\mathbf{Semigp}$ in which we can use the above results to prove some or all of the semigroups $\mathbb{N}\!+\!k$ prime. These categories cannot, of course, contain the semigroups $\textup{Null}(\kappa)$ that were used to prove non -primeness in §2. Here are some restrictions that suffice.
@@ 4. Primeness results, at last
So suppose that for some $s\in S,$ $t,\,t^{\prime}\in T$ we had $\nu_{t}(s)\neq\nu_{t^{\prime}}(s).$ We now invoke weak cancellativity. By left-right symmetry with respect to order of operations, we may assume without loss of generality that there is some $s^{\prime}\in S$ such that $\nu_{t}(s)\,s^{\prime}\neq\nu_{t^{\prime}}(s)\,s^{\prime}.$ But by (3.19), both sides equal $\nu(s\,s^{\prime}),$ a contradiction.
@@
The context of that result concerns categories of algebras whose operations include a binary operation denoted ++ (but not required to be commutative or even associative) and a zeroary operation $0,$ such that $\{0\}$ is a subalgebra, and the identity $0+x=x=x+0$ holds [9, start of §5.4]. The result also involves the “center” of such an algebra, a complicated concept developed in [9, §5.5]; but all we will need to know about it below is that for all elements $b$ in the center, there exists an element $c$ such that $b+c=0$ [9, Definitions 5.11 and 5.12]. (Warning: the word “center” is used in different, though overlapping ways in other chapters of [9].) The result is
@@
Note that primeness in such a category implies primeness in any subvariety, and that a monoid with no nontrivial invertible elements (or more generally, with only finitely many) necessarily satisfies the “finite center” condition of (4.5).
Some examples in the variety of monoids that (4.5) shows to be prime include: (i) all nontrivial submonoids of $\mathbb{N},$ (ii) all submonoids $S$ of $\mathbb{N}\times\mathbb{N}$ such that the submonoids $S\cap(\mathbb{N}\times\{0\})$ and $S\cap(\{0\}\times\mathbb{N})$ are both nontrivial, and $S$ is strictly larger than their sum (e.g., $\{(a,\,b)\in\mathbb{N}\times\mathbb{N}\ |\ a\equiv b\ (\mathrm{mod}~n)\}$ for any $n>1;$ or $\mathbb{N}\times\mathbb{N}\setminus\{(1,0)\}),$ (iii) all submonoids $S$ of $\mathbb{N}\times\mathbb{N}$ such that $S\cap(\mathbb{N}\times\{0\})=\{(0,0)\},$ while $S\cap(\mathbb{N}\times\{1\})$ and $S\cap(\{0\}\times\mathbb{N})$ each have more than one element, and (iv) all free monoids on more than one (possibly infinitely many) generators. (That none of these examples is a nontrivial directly product takes a bit of thought in case (ii) and more in case (iii). I leave these as exercises for the reader.)
@@
On the other hand, we shall see in Example 6.2 below that the additive group of integers is not prime (in the language of that section, “not Tarski-prime”) in the category of groups, from which it easily follows that it is also non-prime in the categories of semigroups and monoids considered above.
5. Two other questions from [9]
@@
An algebra of type <1,1><\!1,1\!> can be thought of as a set given with an action on it of the free monoid on two generators; so in that way the question has a somewhat similar flavor to those studied in the preceding sections.
@@ 6. Some results and examples for groups, comparing
Tarski-primeness and the two sorts of Rhodes-primeness
@@
(iv)  we shall call $G$ Rhodes-prime with respect to direct products if, whenever $G$ is a subquotient of a group in the category in question, and the latter group can be written as a direct product $G_{0}\times G_{1},$ then $G$ is a subquotient of one of $G_{0}$ or $G_{1}$ (condition introduced at [11, p. 482, line 6]); and similarly,
@@
(What we name a subquotient of $G$ in (ii) above was earlier also called a factor of $G,$ e.g., [10, p. 1 et seq.], is called a divisor of $G$ in [11, Def. 1.2.25, p. 32], and is now often called a section of $G$ by group theorists. But “subquotient” is also fairly common.)
We will also follow the common conventions of generally writing $e$ for the identity elements of groups, though for particular groups we may use notation appropriate for them, e.g., 0 for the identity element of the additive group $\mathbb{Z};$ and of calling a group $G$ nontrivial if it has more than one element. (We use “nontrivial” in that sense in the line introducing (iii)-(v) above. In [11] the trivial semigroup is counted as prime with respect to direct and semidirect products, but here we have excluded that group to make the comparison between the Rhodes-primeness conditions and that of Tarski-primeness more straightforward.)
@@
Note that neither of the groups $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ and $\mathbb{Z}_{q}\rtimes\mathbb{Z}$ has a direct product decomposition with a factor isomorphic to $\mathbb{Z}.$ E.g., if $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ did, then a generator of that factor would be a central element of infinite order, hence would have, under the given description of the group, the form (0,ma)(0,\,ma) $(m\neq 0).$ But that would make it an a\!a\!-th power in the whole group, while looking at the $\mathbb{Z}\!$-component in such a direct product decomposition, we see that a generator thereof can’t be an a\!a\!-th power for any $a>1.$ (The groups $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ and $\mathbb{Z}_{q}\rtimes\mathbb{Z}$ are in fact directly indecomposable, but the above argument gives as much as we need.)
On the other hand, suppose we rewrite $G$ as $(\mathbb{Z}_{p}\times\mathbb{Z}_{q})\rtimes(\mathbb{Z}\times\mathbb{Z}),$ using the action of $\mathbb{Z}\times\mathbb{Z}$ under which the first factor $\mathbb{Z}$ affects only the first factor of $\mathbb{Z}_{p}\times\mathbb{Z}_{q},$ and the second affects only the second.
Since $a$ and $b$ are relatively prime, the subgroup of $\mathbb{Z}\times\mathbb{Z}$ generated by the element $(a,b)$ is a direct factor thereof. Explicitly, we can find integers $c$ and $d$ such that $bc+ad=1,$ and use the right action of the invertible matrix $(\begin{smallmatrix}\ b&d\\ \!\!-a&c\end{smallmatrix})$ to get an isomorphism $\mathbb{Z}\times\mathbb{Z}\cong\mathbb{Z}\times\mathbb{Z}$ such that the image of $(a,b)$ generates the second factor. (That matrix or its inverse necessarily has some negative entries, which can be thought of as the way the fact that $\mathbb{Z}$ includes negative integers comes in.) We see from (6.2) that under the action on $\mathbb{Z}_{p}\times\mathbb{Z}_{q},$ that second factor acts trivially.
@@
For finite groups, the relationship between these primeness conditions is quite different. A key result is the Krull-Schmidt Theorem [5, Theorem 3.8, p. 86], which says that every group with ACC and DCC on normal subgroups has a factorization as a direct product of finitely many nontrivial groups which are not themselves nontrivial direct products, and that the factors in such a decomposition are unique up to rearrangement and isomorphism. Thus, every group with those chain conditions which is not a nontrivial direct product is Tarski-prime in the category of such groups. In particular, every nontrivial finite group which is not a nontrivial direct product is Tarski-prime in the category of finite groups. (For a result about finite algebras of much more general sorts than groups, of which this is a special case, see (4.5) above, together with the paragraph that precedes that display, and the second paragraph following it.)
Recall that if a group $G$ has a unique minimal nontrivial normal subgroup $M,$ then $G$ is called monolithic, and $M$ is called the monolith of $G$ [10, sentence beginning at bottom of p. 146]. We can now prove
@@
Though $N$ is a minimal nontrivial normal subgroup of $G,$ it need not be simple, but it will be characteristically simple, i.e., it will have no proper nontrivial subgroup invariant under all its automorphisms (since these include conjugation by all members of $G).$ Hence by [12, 3.3.15, p. 87, and sentence at end of proof, p. 88], $N$ is a direct product of (mutually isomorphic) simple groups. Hence $N^{|G/N|}$ is a direct product of those same simple groups, with possibly more repetitions.
@@
Given a finite monolithic group $G$ that is not Rhodes-prime with respect to direct products, we shall show that neither (i) nor (ii) can hold. The proof that (iii) implies Rhodes-primeness will be more straightforward.
To get the first two results, suppose $G$ is not Rhodes-prime with respect to direct products, i.e., can be written as a homomorphic image $f(H)$ of a subgroup $H$ of a direct product $G_{0}\times G_{1}$ of finite groups, such that $G$ is not a subquotient of $G_{0}$ or of $G_{1}.$ We can clearly replace $G_{0}$ and $G_{1}$ by the subgroups given by the projections of $H$ onto those two factors, so assume that those projections are surjective; i.e., that $H$ is a subdirect product of $G_{0}$ and $G_{1}.$ Let us further assume that, for the given group $G,$ the groups $G_{0},$ $G_{1},$ and $H$ are chosen so as to minimize the order of $H.$
@@
Now let $M_{0}\times\{e\}$ be a nontrivial subgroup of $H_{0}\times\{e\}$ minimal for being normal in $H.$ Since $M_{0}\subseteq H_{0},$ (6.5) tells us that $M_{0}\times\{e\}$ maps one-to-one into $G,$ hence its image is a minimal normal subgroup of $G;$ hence that image is $M.$ Combining with the corresponding observation about a minimal normal subgroup $\{e\}\times M_{1}$ of $\{e\}\times H_{1},$ we get
@@
To get the next assertion, suppose as in (ii) that $G$ is a semidirect product $M\rtimes K$ for some $K<G.$ Since $K$ has trivial intersection with the monolith $M$ of $G,$
@@
Returning to what we proved earlier about $H,$ note that if $f^{-1}(K)<H$ had nontrivial intersection with $H_{0},$ then this would centralize $H_{1},$ hence its image in $G$ would be a subgroup of $K,$ nontrivial by (6.5), that centralized the image of $H_{1},$ which we saw in (6.6) contains $M,$ contradicting (6.8). So $f^{-1}(K)$ has trivial intersection with $H_{0};$ and similarly with $H_{1}.$ In view of (6.4), this forces $f^{-1}(K)$ to be the graph of an isomorphism between subgroups $K_{0}<G_{0}$ and $K_{1}<G_{1},$ each isomorphic to $K,$ and we see that these act on $M_{0}$ and $M_{1}$ as $K$ acts on $M.$ Hence the subgroup $M_{0}\,K_{0}<G_{0}$ (and likewise $M_{1}\,K_{1}<G_{1})$ is isomorphic to $MK=G,$ again contradicting our assumption that $G$ is not a subquotient of $G_{0}$ or $G_{1}.$ This gives the desired contradiction to (ii) for $G$ not Rhodes-prime with respect to direct products.
To prove Rhodes-primeness with respect to direct products in case (iii), suppose a subgroup $H$ of a finite group $G_{0}\times G_{1}$ can be mapped surjectively to $G=\mathbb{Z}_{p^{n}}.$ Then an element mapping to a generator of $G$ must have order divisible by $p^{n}.$ Since the order of an element of $G_{0}\times G_{1}$ is the least common multiple of the orders of its components, one of those components must have order divisible by $p^{n},$ hence some power of that element will have order exactly $p^{n},$ hence the subgroup of $G_{0}$ or $G_{1}$ that it generates will be a subquotient of $G_{0}$ or $G_{1}$ isomorphic to $\mathbb{Z}_{p^{n}}=G.$